# Direct Unlearning Optimization for Robust and Safe Text-to-Image Models

### v1

@@ 4.3 Ablation study
Visual feature unlearning.
The Figure 14 demonstrates the results when removing nudity using only the image compared to using both the image and a prompt. The red teaming was achieved through concept inversion. It is observed that when conducting unlearning using the prompt, DSR does not exceed 0.6. This shows vulnerability to adversarial prompts when utilizing prompt information.
@@ 5 Conclusion and limitation
Figure 14: Ablation study on visual feature unlearning (Concept Inversion). Using prompts for unlearning makes it vulnerable to adversarial prompts.
@@ Appendix A Ablation study
KL-constrained optimization
In this section, we study the effect of KL-constrained optimization. When $\beta\rightarrow 0$, the sigmoid function can be approximated as a linear function, then DUO approaches a naïve combination of gradient ascent for preferred samples and gradient descent for de-preferred samples [65]. As shown in Figure 15, when $\beta\rightarrow 0$, the images are somewhat degraded, and the generation results for unrelated concepts also change. On the other hand, when KL divergence regularization is present, we can confirm that the generation results for unrelated concepts hardly change at all.
@@
Appendix B Implementation details
B.1 Dataset generation
@@
Since $\beta$ linearly multiplies the learning rate at the time near initialization, we rescale the learning rate, dividing it by the same amount each time we increase $\beta$. Using $\beta=100$ as a baseline, we use a learning rate of 1e-4 for nudity and 2e-5 for violence. We use a LoRA rank of 32 for nudity and 2048 for violence. When conducting violence unlearning, we generally observe that the training is more unstable compare to nudity, and we mitigate this by increasing the LoRA rank and lowering the learning rate. Exploring the relationship between dataset characteristics and optimization stability in DUO is an interesting avenue for future work.
@@
For Textual Inversion training, we used the Adam optimizer with a learning rate of lr=5.0e-03, batch size 4, and 3000 gradient steps. We used the same hyperparameters for all unlearning models.
Appendix C Derivation
@@
In (15b), the intractable $p_{\theta}(x_{1:T}^{+}|x_{0}^{+})$ and $p_{\theta}(x_{1:T}^{-}|x_{0}^{-})$ are approximated by the diffusion forward process $q(x_{1:T}|x_{0})$. Now, let’s apply the product rule, i.e., $p(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p(x_{t-1}\mid x_{t})$, to $p_{\theta}(x_{0:T})$ and $p_{\phi}(x_{0:T})$.
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\log{\frac{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}{p_{\phi}(x_{t-1}^{+}|x_{t})}}-\log{\frac{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}{p_{\phi}(x_{t-1}^{-}|x_{t})}}]{)}\right]$ (16a)
$=-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{1:T}^{+}|x_{0}^{+}),q(x_{1:T}^{-}|x_{0}^{-})}$
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}(\log{\frac{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}{q(x_{t-1}^{+}|x_{t})}}-\log{\frac{p_{\phi}^{*}(x_{t-1}^{+}|x_{t})}{q(x_{t-1}^{+}|x_{t})}})-(\log{\frac{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}{q(x_{t-1}^{-}|x_{t})}}-\log{\frac{p_{\phi}^{*}(x_{t-1}^{-}|x_{t})}{q(x_{t-1}^{-}|x_{t})}})]{)}\right]$ (16b)
@@
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\bigg{(}\begin{aligned} &\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{+}|x_{t})\big{)}-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{+}|x_{t})\big{)}\\ &-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{-}|x_{t})\big{)}-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{-}|x_{t})\big{)}\end{aligned}\bigg{)}]{)}\right]$
Finally, by applying the variational lower bound of the diffusion loss, we obtain the upper bound of the above equation as follows [23]:
@@ 18. Answer: [Yes]
  19. Justification: In the Appendix B, we provide every details to reproduce our results.
@@ 28. Answer: [Yes]
  29. Justification: We describe all the hyperparameters, optimizer, and dataset in Appendix B.
@@ 38. Answer: [Yes]
  39. Justification: We report computation resources and time complexity of our algorithm in Appendix B.

### v2

@@ Appendix A Ablation study
KL-constrained optimization.
In this section, we study the effect of KL-constrained optimization. When $\beta\rightarrow 0$, the sigmoid function can be approximated as a linear function, then DUO approaches a naïve combination of gradient ascent for preferred samples and gradient descent for de-preferred samples [66]. As shown in Figure 14, when $\beta\rightarrow 0$, the images are somewhat degraded, and the generation results for unrelated concepts also change.
@@
Figure 15: Ablation study on the number of synthesized image pairs. Different colors represent Ring-A-Bell results for nudity unlearning, varying the number of image pairs from 1 to 1024.
The number of synthesized image pairs
Figure 15 demonstrates how varying the number of synthesized image pairs affects the Pareto curve on the Ring-A-Bell nudity benchmark. The figure clearly shows that when the number of pairs is less than 64, there is a noticeable improvement in the Defense Success Rate. However, increasing the number beyond 64 pairs does not yield significant changes in the Pareto curve. This analysis suggests that 64 pairs provide a good balance between performance and computational efficiency for our method.
Reproducibility
We have conducted experiments using different random seeds (1 to 4) to demonstrate the stability and reproducibility of our results. Figure 17 shows that DUO’s Pareto curve is not sensitive to random seed variation and consistently outperforms baseline methods. Figure 18 provides qualitative evidence of similar unlearning results across different seeds.
Figure 16: Quantitative result of UnlearnDiffAtk on nudity.
Appendix B Additional experiments
UnlearnDiffAtk [68]
We conducte additional experiments using UnlearnDiffAtk [68], a state-of-the-art white-box attack method designed for assessing the robustness of unlearned diffusion models.
As illustrated in Figure 16, DUO achieves Pareto-optimal performance compared to existing baselines when subjected to UnlearnDiffAtk. This means that DUO provides the best trade-off between maintaining model performance and resisting attacks, outperforming other methods in both aspects simultaneously.
Figure 17: Quantitative result on various seed numbers. DUO (Reproduce) shows the experimental results using different random seeds. DUO consistently outperforms other methods, maintaining its superior performance across different seed numbers, demonstrating its stability and effectiveness.
Figure 18: Qualitative result on various seed numbers. We used $\beta=500$ for both nudity and violence. We use Refer to caption and blurring for publication purposes.
Impact of DUO on Unrelated Features
To further validate prior preservation performance of our approach, we evaluated the model’s ability to generate visually similar but safe concepts. For example, we compare the generation of red images (e.g., ketchup, strawberry jam) after removing the Violence concept, which is closely related to "Blood". Table 3 below presents mean and standard devidation of LPIPS scores between 128 images generated from the unlearned model and the pretrained model. Lower scores indicate less impact on unrelated features. These results demonstrate that DUO effectively maintains the capability to generate visually similar but unrelated concepts compared to the existing methods. Figure 19 shows qualitative results of prior preservation for visually similar concepts.
Table 3: Impact of unlearning on visually similar concept generation. We report the mean $\pm$ standard deviation of LPIPS scores between 128 images generated from the unlearned model and those from the pretrained model.
Unlearned Concept Safe Concept ESD UCE SPM DUO (black box) DUO (white box)
Nudity Woman $0.58\pm 0.11$ $0.42\pm 0.16$ $0.31\pm 0.15$ $0.33\pm 0.12$ $0.55\pm 0.17$
Man $0.58\pm 0.10$ $0.31\pm 0.17$ $0.15\pm 0.13$ $0.14\pm 0.09$ $0.20\pm 0.13$
Violence Ketchup $0.69\pm 0.15$ $0.51\pm 0.16$ $0.20\pm 0.15$ $0.23\pm 0.12$ $0.35\pm 0.14$
Tomato sauce $0.58\pm 0.19$ $0.38\pm 0.16$ $0.11\pm 0.13$ $0.18\pm 0.12$ $0.28\pm 0.12$
Strawberry jam $0.56\pm 0.13$ $0.42\pm 0.15$ $0.13\pm 0.12$ $0.20\pm 0.11$ $0.31\pm 0.12$
Figure 19: Quantitative result on generating close concept related to removed concept. (a) Results of generating "man" and "woman" after removing nudity. (b) Results of generating "ketchup", "strawberry jam" and "tomato sauce" after removing violence.
Stable-Diffusion 3
In Figures 20, and 21, we successfully applied DUO to Stable Diffusion 3 (SD3), which uses the transformer-based mmDiT architecture [10]. The evaluation process for SD3 was identical to that used for SD1.4v. In the Pareto curve, each point from left to right represents $\beta\in\{100,250,500,1000,2000\}$, with a consistent learning rate of 3e-5 across all experiments.
Figure 20: Quantitative result of SD3 on nudity. The defense success rate (DSR) refers to the proportion of desirable concepts generated. Prior preservation represents the perceptual similarity of images generated by the prior model and the unlearned model.
Figure 21: Qualitative result of SD3 on nudity. DUO effectively removes nudity while preserving the model’s ability to generate unrelated concepts. We use Refer to caption for publication purpose.
Additionally, Table 4 shows FID scores on MS COCO. It demonstrates that applying DUO to SD3 maintains comparable fidelity to the pretrained model.
Table 4: FID for SD3 unlearned nudity.
Method FID ($\downarrow$)
SD3 21.83
DUO ($\beta=500$) 21.26
DUO ($\beta=250$) 20.49
Appendix C Implementation details
C.1 Dataset generation
@@
Since $\beta$ linearly multiplies the learning rate at the time near initialization, we rescale the learning rate, dividing it by the same amount each time we increase $\beta$. Using $\beta=100$ as a baseline, we use a learning rate of $43\times 10^{-}$ 4, the batch size of 4, and LoRA rank of 32. When conducting violence unlearning, we generally observe that the training is more unstable compared to nudity, and we mitigate this by training LoRA to unlearn each concept and combine them as violence LoRA. Exploring the relationship between dataset characteristics and optimization stability in DUO is an interesting avenue for future work.
@@
For Textual Inversion training, we used the Adam optimizer with a learning rate of $35\times 10^{-}$ 3, batch size 4, and 3000 gradient steps. We used the same hyperparameters for all unlearning models.
Appendix D Derivation
D.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
For completeness, we provide the derivation of the solution of KL-constrained reward maximization here. For more detailed motivation and explanation of the derivation process, we refer readers to the DPO paper [44] Appendix A.1.
Eq. (4) $\rightarrow$ Eq. (5)
Start from Eq. (4).
$\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}$ $[r(x_{0})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0})||p_{\phi}(x_{0})]$
$=\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}\left[r(x_{0})-\beta\log\frac{p_{\theta}(x_{0})}{p_{\phi}(x_{0})}\right]$
$=\min_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}\left[\log\frac{p_{\theta}(x_{0})}{p_{\phi}(x_{0})}-\frac{1}{\beta}r(x_{0})\right]$
$=\min_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}\left[\log\frac{p_{\theta}(x_{0})}{\frac{1}{Z}p_{\phi}(x_{0})\exp\left(\frac{1}{\beta}r(x_{0})\right)}-\log Z\right]$
where we have partition function:
$Z\triangleq\sum_{x_{0}}p_{\phi}(x_{0})\exp{\frac{1}{\beta}r(x_{0})}$
Now define the probability $p^{*}(x_{0})$ as:
$p^{*}(x_{0})\triangleq\frac{1}{Z(x_{0})}p_{\phi}(x_{0})\exp\left(\frac{1}{\beta}r(x_{0})\right)$
which is a valid probability distribution as $p^{*}(x_{0})\geq 0$ for all $x_{0}$ and $\sum_{x_{0}}p(x_{0})=1$. Since $Z$ is not a function of $x$, we can then re-organize the final objective as:
$\min_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0})}\left[\log\frac{p_{\theta}(x_{0})}{p^{*}(x_{0})}-\log Z\right]$
Now, since $Z$ does not depend on $p_{\theta}$, the minimum is achieved by $p_{\theta}(x_{0})=p^{*}(x_{0})$. Hench we have the optimal solution:
$p_{\theta}(x_{0})=p^{*}(x_{0})=p_{\phi}(x_{0})\exp\left(r(x_{0})/\beta\right)/Z$
D.2 Diffusion-DPO
For completeness, we provide the derivation of Diffusion-DPO here. For more detailed motivation and explanation of the derivation process, we refer readers to the Diffusion-DPO paper [57].
@@ D.2 Diffusion-DPO
$\approx-\mathbb{E}_{x_{0}^{+},x_{0}^{-},q(x_{1:T}^{+}|x_{0}^{+}),q(x_{1:T}^{-}|x_{0}^{-})}$
In Eq. (15b), the intractable $p_{\theta}(x_{1:T}^{+}|x_{0}^{+})$ and $p_{\theta}(x_{1:T}^{-}|x_{0}^{-})$ are approximated by the diffusion forward process $q(x_{1:T}|x_{0})$. Now, let us apply the product rule, i.e., $p(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p(x_{t-1}\mid x_{t})$, to $p_{\theta}(x_{0:T})$ and $p_{\phi}(x_{0:T})$.
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\log{\frac{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}{p_{\phi}(x_{t-1}^{+}|x_{t})}}-\log{\frac{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}{p_{\phi}(x_{t-1}^{-}|x_{t})}}]{)}\right]$
@@
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\bigg{(}\begin{aligned} &\log{{p_{\theta}^{*}(x_{t-1}^{+}|x_{t})}/{q(x_{t-1}^{+}|x_{t})}}-\log{{p_{\phi}^{*}(x_{t-1}^{+}|x_{t})}/{q(x_{t-1}^{+}|x_{t})}}\\ &-\log{{p_{\theta}^{*}(x_{t-1}^{-}|x_{t})}/{q(x_{t-1}^{-}|x_{t})}}+\log{{p_{\phi}^{*}(x_{t-1}^{-}|x_{t})}/{q(x_{t-1}^{-}|x_{t})}}\end{aligned}\bigg{)}]{)}\right]$
@@
$\left[\log\sigma{(}[\beta\sum_{t=1}^{T}\bigg{(}\begin{aligned} &\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{+}|x_{t})\big{)}-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{+}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{+}|x_{t})\big{)}\\ &-\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}^{-}|x_{t})\big{)}+\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}^{-}|x_{t})\big{\|}p_{\phi}^{*}(x_{t-1}^{-}|x_{t})\big{)}\end{aligned}\bigg{)}]{)}\right]$
Finally, let us express the $\mathbb{D}_{\text{KL}}(q||p)$ using the $L_{\text{dsm}}$, i.e., $\|\epsilon-\epsilon_{\theta}\|_{2}^{2}$. According to Eq. (48) from Kingma et al. [24], the following relationship holds:
$\mathbb{E}_{x_{t}\sim q(x_{t}|x_{0})}\left[\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}|x_{t})\big{)}\right]=\mathbb{E}_{\epsilon\sim N(0,I)}w(\lambda_{t})\|\epsilon-\epsilon_{\theta}(x_{t})\|_{2}^{2}$ (17)
where $\lambda_{t}$ is the signal-to-noise ratio (SNR), defined as $\lambda_{t}=\frac{\alpha}{1-\alpha}$, and $w$ is a coefficient factor that depends on the SNR, specifically $w(\lambda_{t})=\frac{1}{\lambda_{t}}\frac{d}{dt}\lambda_{t}$. In practice, when training the diffusion model, $w$ is often treated as a constant [21]. By using Eq. (17) to replace each $\mathbb{D}_{\text{KL}}\big{(}q(x_{t-1}|x_{t})\big{\|}p_{\theta}^{*}(x_{t-1}|x_{t})\big{)}$, we obtain the following:
$L_{\text{Diffusion-DPO}}\leq$ $-\mathbb{E}_{x_{0}^{+},x_{0}^{-},x_{t}^{+}\sim q(x_{t}^{+}|x_{0}^{+}),x_{0}^{-}\sim q(x_{t}^{-}|x_{0}^{-})}$ (18)
@@ 18. Answer: [Yes]
  19. Justification: In Appendix C, we provide every details to reproduce our results.
@@ 28. Answer: [Yes]
  29. Justification: We describe all the hyperparameters, optimizer, and dataset in Appendix C.
@@ 38. Answer: [Yes]
  39. Justification: We report computation resources and time complexity of our algorithm in Appendix C.